OpenStudy (pamstevens):

I have to solve this equation by making the appropriate substitution. Does anyone know how to work this problem? x+3x^1/2-10=0

OpenStudy (sdfgsdfgs):

try substitute y=x^1/2 so x=y^2

OpenStudy (pamstevens):

Ok I see what you are trying to do but does that mean I don't combine the x"s?

OpenStudy (sdfgsdfgs):

nothing to combine there as u cannot combine x and x^1/2 use the substitution n solve for u first

OpenStudy (pamstevens):

Can you show me how you would set up the problem using those substitutions?

OpenStudy (sdfgsdfgs):

y=x^1/2 y^2=x put in the original eqn...